hannahcj:

Nikki can drive her car 22 miles on each gallon of gasoline. If Nikki drives m miles, which expression represents the number of gallons used? m - 22 22m 22/m m/22

jhonyy9:

22 miles on 1 gallon gasoline m miles on x gallons cross multiplie and calcule the x value x = m*1/22 = m/22 hope helped

jhonyy9:

@XioGonz

hannahcj:

Yes it did! Thank you.

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 22 miles on 1 gallon gasoline m miles on x gallons cross multiplie and calcule the x value x = m*1/22 = m/22 hope helped \(\color{#0cbb34}{\text{End of Quote}}\) Correct! Nice job (:

hannahcj:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 22 miles on 1 gallon gasoline m miles on x gallons cross multiplie and calcule the x value x = m*1/22 = m/22 hope helped \(\color{#0cbb34}{\text{End of Quote}}\) I gave you a medal.

jhonyy9:

ty

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 22 miles on 1 gallon gasoline m miles on x gallons cross multiplie and calcule the x value x = m*1/22 = m/22 hope helped \(\color{#0cbb34}{\text{End of Quote}}\) Correct! Nice job (: \(\color{#0cbb34}{\text{End of Quote}}\) ty

hannahcj:

Your welcome. Have a great day!

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 22 miles on 1 gallon gasoline m miles on x gallons cross multiplie and calcule the x value x = m*1/22 = m/22 hope helped \(\color{#0cbb34}{\text{End of Quote}}\) Correct! Nice job (: \(\color{#0cbb34}{\text{End of Quote}}\) ty \(\color{#0cbb34}{\text{End of Quote}}\) Np!